Using Epsilon and Delta definition, show that limit of x approches to 2 for 3x-5 is equal to 1.

1
Expert's answer
2021-11-08T19:41:16-0500

Using Epsilon and Delta definition, let us show that "\lim\limits_{x\to 2}3x-5 =1."

Let "\varepsilon>0" be arbitrary. Put "\delta=\frac{1}3\varepsilon." Then for any "\varepsilon>0" there exists "\delta>0" such that if "|x-2|<\delta" then "|(3x-5)-1|=|3x-6|=3|x-2|<3\delta=3\frac{1}3\varepsilon=\varepsilon." We conclude that "\lim\limits_{x\to 2}3x-5 =1."